OpenStudy (anonymous):

the acute angle measure of the hands of a clock is (blank) at 5:40. need alot of help plz medal given out

hero (hero):

Okay, well, basically, you have to figure out the measure of the hour hand and the minute hand, then find the difference between both measures.

hero (hero):

So basically, you know that there are 12 evenly spaced segments for the hour hand so: \[\frac{5}{12} = \frac{x}{360}\] You solve for x for the hour hand Then for the minute hand, there are 60 evenly spaced segments so solve for y: \[\frac{40}{60} = \frac{y}{360}\]

OpenStudy (anonymous):

so how would you do that?

hero (hero):

Lastly, you want to find |x - y|

hero (hero):

x and y represents the degree measures of each hand

OpenStudy (anonymous):

then what I keep messing up

hero (hero):

What did you get for the result?

OpenStudy (anonymous):

87

OpenStudy (anonymous):

degrees

hero (hero):

Hmmm? What did you get for x and y?

hero (hero):

Oh, I see...

OpenStudy (anonymous):

i dont get it the way your doing it

hero (hero):

I can't help you with this because there's more to it than just that.

OpenStudy (anonymous):

why

hero (hero):

There are other things you have to account for.

OpenStudy (anonymous):

not really

hero (hero):

Because the hour hand isn't exactly on the five.

hero (hero):

at 5:40

OpenStudy (anonymous):

just the degrees between the hands

hero (hero):

Well, I know there's more to it because it says "acute angle measure". The actual measurement between 5 and 8 is 90 degrees. But the 5 is closer to the 6 when the minute hand gets to the 8.

hero (hero):

I can't help you with this because I don't remember how to do it.

OpenStudy (anonymous):

i know but cant you solve it

hero (hero):

Yes, I could figure it out with time.

OpenStudy (anonymous):

your the smartest on this thing and i need help

OpenStudy (anonymous):

thank you i have alot of time

hero (hero):

I found a cheat way to do it here: http://en.wikipedia.org/wiki/Clock_angle_problem

OpenStudy (anonymous):

you cant move the hands

hero (hero):

The angle for the hour hand would be: \[\frac{1}{2}\left(60(5) + 40\right) = 170\] The angle for the minute hand would be \[6(40) = 240\] And 240 - 170 = 70

hero (hero):

I think...

OpenStudy (anonymous):

so 70 degrees?

hero (hero):

Wait..hang on...

OpenStudy (anonymous):

k

hero (hero):

Yes, that should be it

OpenStudy (anonymous):

alright ill try it

OpenStudy (anonymous):

YOU WERE RIGHT THX MAN